Given a sequence ((xn,yn)) is R2 .prove that if ((xn,yn)) is bounded ,then (xn) and (yn) are bounded.

1
Expert's answer
2022-04-18T00:28:31-0400

ANSWER

By the definition, the sequence "\left\{ \left( { x }_{ n }{ ,y }_{ n } \right) \right\}" is bounded in "R^2" if there exists "M>0" such that

"\left\| \left( { x }_{ n },{ y }_{ n } \right) \right\| =\sqrt { { \left( { x }_{ n } \right) }^{ 2 }+{ \left( { y }_{ n } \right) }^{ 2 } } \le M" for all "n\in N" .


Therefore

"\left| { x }_{ n } \right| =\sqrt { { \left( { x }_{ n } \right) }^{ 2 }\quad } \le \sqrt { { \left( { x }_{ n } \right) }^{ 2 }+{ \left( { y }_{ n } \right) }^{ 2 } } \le M" for all "n\in N" .

and

"\left| { y }_{ n } \right| =\sqrt { { \left( { y }_{ n } \right) }^{ 2 }\quad } \le \sqrt { { \left( { x }_{ n } \right) }^{ 2 }+{ \left( { y }_{ n } \right) }^{ 2 } } \le M"

Equivalent to

"-M\leq x_{n}\leq M" . for all "n\in N"

and

"-M\leq y_{n}\leq M" for all "n\in N"

So, the sequences "\left\{ \left( { x }_{ n } \right) \right\} ," "\left\{ \left( {y }_{ n } \right) \right\}" are bounded in "R" .